# Black Hole Superradiance of Interacting Multi-Field

### v2

@@ III.2 Interactions between bound states and unbound states of $\psi$
Besides radiating $\varphi$, the superradiant mode could also lose energy to spacial infinity by radiating $\psi$ via a process depicted by the right diagram in Fig. 4. This process, hereinafter referred to as the $3{\bar{\psi}}$-process, is described by
@@
which are obtained from Eqs. (13) and (14), with $\psi$ on the r.h.s. being the background ${\bar{\psi}}_{0}$. Here we do not take the non-relativistic ansatz for $\psi$ because the radiated $\psi$ is unbounded and can be relativistic. The effects of $3{\bar{\psi}}$-process on the growth rate of ${\bar{\psi}}_{211}$ can be investigated by estimating the radiation power of $\psi$.
@@
Figure 7: Parameter space of interaction corrections. From left to right, the plots show the regions of nontrivial $\delta\omega_{211}^{sE}$ (in red), $\delta\omega_{211}^{sB}$ (in blue), and $\Gamma_{R}$ (in green). Again, by assuming ${\bar{\psi}}_{211}$ to be the fastest growing mode, we first exclude the white regime in the middle and right plots. For $\delta\omega_{211}^{sE}$ to be non-trivial, it further requires $q\lesssim 2$. Moreover, we distinguish the regions that ${\rm Im}\delta\omega_{211}^{sB}>0$ (in dark blue), ${\rm Im}\delta\omega_{211}^{sB}<0$ (in light blue), and ${\rm Im}\delta\omega_{211}^{sB}$ changes from positive to negative as the black hole spin evolves from $\sim 1$ to $a_{crit}$ (in blue).
Again, $\varphi$ can be solved from Eq. (34) with the Green function method,
@@
Figure 8: Parameter space of the dominating interaction corrections. We show the regions that the term of $\gamma_{E}$ (in red), the $\gamma_{B}$ (in blue) and $\gamma_{R}$ (in green) might dominate the superradiant evolution. We further exclude the region that $\epsilon_{eq}>\epsilon_{th}\equiv N_{th}/GM^{2}$, cf. Fig. 7, where the perturbative treatment may break down. Moreover, we also show the contours of $\tilde{\lambda}_{c}$ in each color regions. We expect that the interaction correction can interrupte the superradiant growth of $\psi_{211}$ significantly, and lead to the quasi-equilibrium stage if $\tilde{\lambda}>\tilde{\lambda}_{c}$. Given the value of $\tilde{\lambda}_{c}$, we find that the superradiant growth of ${\bar{\psi}}_{211}$ can be easily affected by $\varphi$ even with very weak coupling, i.e., $\lambda/\mu\ll 1$.
@@ IV Superradiance Evolution of interacting fields
In principle, the black hole mass also evolves during the superradiance evolution. Nevertheless, we treat the black hole mass to be constant, since the black hole mass typically varies by a small amount during superradiance. This is especially the case in the presence of the interaction, as the cloud tends to grow to a smaller occupation number. With this treatment, $\alpha$ can also be considered as a constant. We could also take into account gravitational wave radiations by appending $\gamma_{GW}\epsilon_{211}^{2}$ on the r.h.s. of both Eq. (44) with $\gamma_{GW}=\alpha^{14}/160$ [65, 66]. While we are interesting in the case with $\epsilon_{eq}$ derived from the interaction corrections is less than 1, the gravitational wave radiations leads to an effective $\epsilon_{eq}\gg 1$ due to the weak nature of gravitational interactions, and therefore is irrelevant for the discussion. Fig. 9 shows a typical evolution in the presence of interaction, which is obtained by solving Eqs. (44) and (45) numerically. We find that the superrdiant mode typically experiences three stages: exponential growth, quasi-equilibrium, and power-law decay, which can be understood as follows.
Initially, $\epsilon_{211}$ is very small, and its evolution is dominated by the linear term in Eq. (44), in which case the mode will grow exponentially. The exponential growth continues until the $\gamma_{211}$ becomes almost zero due to black hole spin down in which case the cloud saturated, or the interaction terms in Eq. (44) becomes significant. In the former case, the later evolution would be similar to that of free fields. In the later case, the exponential growth could be interrupted by the interaction corrections, and the superradiant cloud researches a quasi-equilibrium stage. By requiring the l.h.s. of Eq. (44) to be zero, one can define the dimensionless occupation number at the quasi-equilibrium stage $\epsilon_{eq}$. We expect the cloud reaches the quasi-equilibrium stage before saturation if $\epsilon_{eq}<\epsilon_{max}$.
Although different processes may take place simultaneously as shown in Fig. 7, typically only one process dominates the correction, which allowing us to simplify the equation by neglecting the sub-dominating correction. In Fig. 8, we show the parameter space of the dominating process, with the contours denoting the critical $\tilde{\lambda}_{c}$ defined as $\epsilon_{eq}(\tilde{\lambda}_{c})=\epsilon_{max}$. Namely, the cloud would reach the quasi-equilibrium stage before it saturates as in the single field case if $\tilde{\lambda}>\tilde{\lambda}_{c}$. Here $\tilde{\lambda}$ relates to $\lambda$ via
@@
Given the value of $\tilde{\lambda}_{c}$, we find that the cloud reaches the equilibrium stage with very tiny $\epsilon_{eq}$ even if $\lambda/\mu\ll 1$. Depending on the dominating process, the subsequent evolution may be slightly different, and will be discussed as follows.
@@ IV.1 Case A
Without the interruption of other superradiant modes, the quasi-equilibrium stage continues until $\epsilon_{eq}$ approaches to 0 as the black hole spin approaches to the critical spin ${\tilde{a}}_{crit}$. While Eq. (50) can be solved analytically, we shall not write down the lengthy expression of $a$, but are rather interested in the evolution when $\tilde{a}$} is close to ${\tilde{a}}_{crit}$, in which case
@@ IV.2 Case B
with $\epsilon_{eq}\simeq(\gamma_{211}/\gamma_{R})^{1/2}$. While $\epsilon_{211}$ could also reaches $\epsilon_{eq}$ as in case A, the equilibrium may not be always stable. Considering a small deviation $\delta\epsilon$ from $\epsilon_{eq}$ and by perturbing Eq. (53), we have
@@
Comparing the two time scales $\tau_{var}$ and $\tau_{re}$ leads to a threshold in term of black hole spin
@@
below which the quasi-equilibrium stage becomes unstable. If the quasi-equilibrium is stable as $a$ approaches to $a_{crit}$, the cloud decays with $\epsilon_{211}\propto t^{-1}$. Otherwise, $\epsilon^{\prime}_{211}$ is dominated by the $\gamma_{R}\epsilon_{211}^{3}$ term, and the cloud decays with $\epsilon_{211}\propto t^{-1/2}$.
@@ IV.3 Case C
In region C of Fig. 8, the $\gamma_{B}$ term typically first dominates the interaction correction. The evolution is similar to that in case A, except that the black hole spin stops decreasing once reaching the quasi-equilibrium stage. The black hole spin and the occupation number remain almost constant until the $\gamma_{R}$ term becomes significant and takes over the evolution.
@@ IV.4 Cloud Collapse
and investigating how $R$ deviates from the Bohr radius $r_{B}$ in the presence of the effective self-interaction. In the non-relativistic limit, the action of $\bar{\psi}$} reduces to
@@
with $\rho_{\rm BH}=M\delta^{3}({\bf r})$. $\varphi$ is a mediator. Taking into account the contribution of s, t and u channels, we have
@@
where $\tilde{R}=R/r_{B}$ and $I_{p}$ is a dimensionless factor containing contribution of $\ell m=22,20,00$, with its numerical values of each component shown in Fig. 11. The four terms in Eq. (61) correspond to kinetic energy, black hole gravity, quartic interaction and cloud self-gravity respectively. The extrema of $V_{\rm eff}$ locates at
@@ V Implications on Observations
Even with a sufficient growth of the cloud, the existence of the other fields can change the evolution of the cloud, including its maximum mass, decay rate and lifetime, and hence leads to different gravitational wave signals comparing to the single field case. These effects should be taken into account in when searching for dark particles with black hole superradiance.
@@ VI Conclusion and Discussion
In this work, we investigated black hole superradiance of interacting multiple fields. Taking a two-field model for demonstration, we discuss the possible interactions between the bound and unbound states of the two fields, and their impacts on the superradiant growth via a perturbative approach. We further analyze the superradiant growth of the most fastest mode in the presence of the interactions. We find that the superradiant mode typically experiences three stages: the exponential growth stage, the quasi-equilibrium stage and the power-law decay stage. We also find that the superradiance is typically suppressed if the superradiant field couples to another field even with a very weak coupling strength.
While we taking a two-field model for demonstration, inclusion of more fields should not alter the picture essentially. This is especially the case when there is a large hierarchy in the field mass such that one of the fields grows efficiently, and another field contributes most to the interaction corrections. Nevertheless, one could consider the superradiant growth of the other modes, for example, ${\bar{\psi}}_{322}$, ${\bar{\varphi}}_{211}$ or ${\bar{\varphi}}_{322}$ in our setup, which may play a role in the late evolution of ${\bar{\psi}}_{211}$ in certain parameter regions. Including such modes into the discussion is beyond the scope of this work, and will be studied in future work. It is also interesting to extend the discussion to the spin-1/2 and spin-1 fields, and to consider other possible interactions and their effects on superradiant growth.
It is possible that the dark sector consists of many species of interacting dark particles, rather than just one species of dark particle. As we demonstrated in this work that the existence of the other fields can have important impacts on the observational signatures of superradiance, the superradiance based constraints on dark particles derived from single-field analyses should be revised in the presence of interactions. The observational consequences of multi-field black hole superradiance, such as black hole spin distribution, gravitational wave emission from the clouds, and other potential signals of superradiant cloud, are also interesting questions to be studied in future.
@@ Appendix A Ultralight scalar fields in the weak gravity limit of the Kerr black hole background
with $\hat{L}^{2}$ being the total angular momentum operator. In the second line of above equation, we have used the equation of free $\psi$ at zeroth order of $GM/r$ to eliminate the $\partial_{t}^{2}$ operator. For non-relativistic states, the wavenumber $k$ is much less than $\mu$. Therefore, $\hat{K}^{2}\sim k^{2}$ can be neglected, and we have
@@
for non-relativistic states in the weak gravity limit.
@@
The general solution to Eq. (70) are well known in the literature. By imposing the regular boundary condition, one can select a set of relevant eigenstates $\{u_{n\ell m}\}$ and $\{u_{k\ell m}\}$, which satisfy
@@
In the literature, the bound states are also often described with $\bar{\psi}$} defined by the non-relativistic ansatz (6). By substituting ansatz (6) into Eq. (67) and keeping leading order terms in $\alpha$, we can find that $\bar{\psi}$} satisfies
@@
with $\xi_{n\ell m}\simeq\omega_{n\ell m}-\mu$. For relativistic states, one should take $\hat{K}^{2}$ into account. Nevertheless, when we calculation the radiation power of the $3{\bar{\psi}}$-process, we take the flat-space approximation for simple, in which case the eigenfunction of the unbound states can be obtained by taking $u_{k\ell m}$ to the limit of $kr_{B}\rightarrow\infty$, and the radial function $R_{k\ell}$ reduces to spherical Bessel function.
Appendix B Green function method
In this appendix, we shall show how to solve the inhomogeneous Klein-Gordon equation [^5] Here we use the $\psi$ field to demonstrate the method, while in the main text, the Green function method is applied to solve Eq. (17) and Eq. (34) for the $\varphi$ field.
@@ Appendix B Green function method
on the Kerr background with the Green function method. We shall set the Newton’s constant $G=1$ in this appendix to avoid confusion with the Green function. In the Boyer-Lindquist coordinates, the equation is separable with
@@
B.1 The radial Green function
The radial function can be solved with the radial Green function,
@@ B.1 The radial Green function
where $r_{*}$ is the tortoise coordinate. The Green function can be construct as
@@
which describes the in-going wave at infinity. Since there are only two independent solutions, the above solutions are related by
@@ B.2 Relation to the eigenstates
Figure 13: Analyticity of $\zeta(z)$ and the integral contour in integral (87). The branch cut is chosen at $(-\infty,-\mu]$ and $\mu,\infty)$. The left plot shows the principal Riemann sheet, and the right plot shows the second Riemann sheet. There are poles at $\pm\Omega$ (the red dots), and at the zeros of Wronskian (the black dots). The bound state poles (${\rm Re}\omega<\mu$) are on the principle Riemann sheet, while the qausi-normal mode poles are on the second Riemann sheet.
Now we show that Eq. (80) can be expanded with the eigenfunctions of homogeneous radial equation, i.e. Eq. (B) with $I=0$. We shall show this by constructing an integral
@@
Recall that $\Omega$ is the forced oscillation frequency of the source term $s(x)$. The integration contour and the analyticity of $\zeta(z)$ are shown in Fig. 13. Due to the multi-value of $k=\pm\sqrt{z^{2}-\mu^{2}}$, $\zeta(z)$ has two branch point $z=\pm\mu$ and we define the integral in the Riemann sheet of ${\rm Im}\,k>0$ to fit the out-going boudnary condition, which means the branch cut is $(-\infty,-\mu]$ and $\mu,\infty)$.
@@
such that it approaches to the normalized eigenfunctions discussed in App. A in the weak gravity limit.
@@
For $s(x)=\tilde{s}(\mathbf{r})e^{-i\Omega t}+c.c.$ with general angular dependence, one can consider the full Green function [68]
@@
and use Eq. (93) when computing the radial part. Specifically, in the weak gravity limit, we have $A_{\mathrm{in}}A_{\mathrm{out}}=1$, $f_{n}(r)S_{\omega\ell m}(\theta)e^{im\phi}$ and $f_{H}(r)S_{\omega\ell m}(\theta)e^{im\phi}$ reduce to $u_{n\ell m}(\mathbf{r})$ and $u_{k\ell m}(\mathbf{r})$ discussed in App. A respectively. Therefore, the solution to Eq. (19) can be given by Eq. (LABEL:eq:expPhi).

### v3

@@ III.2 Interactions between bound states and unbound states of $\psi$
Besides radiating $\varphi$, the superradiant mode could also lose energy to spacial infinity by radiating $\psi$ via a process depicted by the right diagram in Fig. 6. This process, hereinafter referred to as the $3{\bar{\psi}}$ process, is described by
@@
which are obtained from Eqs. (13) and (14), with $\psi$ on the rhs being the background ${\bar{\psi}}_{0}$. Here we do not take the nonrelativistic ansatz for $\psi$ because the radiated $\psi$ is unbounded and can be relativistic. The effects of the $3{\bar{\psi}}$ process on the growth rate of ${\bar{\psi}}_{211}$ can be investigated by estimating the radiation power of $\psi$.
@@
Figure 7: Parameter space of interaction corrections. From left to right, the plots show the regions of nontrivial $\delta\omega_{211}^{sE}$ (in red), $\delta\omega_{211}^{sB}$ (in blue), and $\Gamma_{R}$ (in green). Again, by assuming ${\bar{\psi}}_{211}$ to be the fastest growing mode, we first exclude the white regime in the middle and right plots. For $\delta\omega_{211}^{sE}$ to be nontrivial, it further requires $q\lesssim 2$. Moreover, we distinguish the regions that ${\rm Im}\delta\omega_{211}^{sB}>0$ (in dark blue), ${\rm Im}\delta\omega_{211}^{sB}<0$ (in light blue), and ${\rm Im}\delta\omega_{211}^{sB}$ changes from positive to negative as the black hole spin evolves from $\sim 1$ to $a_{crit}$ (in blue).
Again, $\varphi$ can be solved from Eq. (34) with the Green’s function method,
@@
Figure 8: Parameter space of the dominating interaction corrections. We show the regions that the term of $\gamma_{E}$ (in red), $\gamma_{B}$ (in blue), and $\gamma_{R}$ (in green) might dominate the superradiant evolution. We further exclude the region that $\epsilon_{eq}>\epsilon_{th}\equiv N_{th}/GM^{2}$, cf. Fig. 7, where the perturbative treatment may break down. Moreover, we also show the contours of $\tilde{\lambda}_{c}$ in each color regions. We expect that the interaction correction can interrupt the superradiant growth of $\psi_{211}$ significantly, and lead to the quasiequilibrium stage if $\tilde{\lambda}>\tilde{\lambda}_{c}$. Given the value of $\tilde{\lambda}_{c}$, we find that the superradiant growth of ${\bar{\psi}}_{211}$ can be easily affected by $\varphi$ even with very weak coupling, i.e., $\lambda/\mu\ll 1$.
@@ IV Superradiance Evolution of interacting fields
In principle, the black hole mass also evolves during the superradiance evolution. Nevertheless, we treat the black hole mass to be constant, since the black hole mass typically varies by a small amount during superradiance. This is especially the case in the presence of the interaction, as the cloud tends to grow to a smaller occupation number. With this treatment, $\alpha$ can also be considered as a constant. We could also take into account gravitational wave radiations by appending $\gamma_{GW}\epsilon_{211}^{2}$ on the rhs of both Eq. (44) with $\gamma_{GW}=\alpha^{14}/160$ [16, 63]. While we are interested in the case where $\epsilon_{eq}$ derived from the interaction corrections is less than 1, the gravitational wave radiations lead to an effective $\epsilon_{eq}\gg 1$ due to the weak nature of gravitational interactions, and therefore is irrelevant for the discussion. Figure. 9 shows a typical evolution in the presence of interaction, which is obtained by solving Eqs. (44) and (45) numerically. We find that the superradiant mode typically experiences three stages: exponential growth, quasiequilibrium, and power-law decay, which can be understood as follows.
Initially, $\epsilon_{211}$ is very small, and its evolution is dominated by the linear term in Eq. (44), in which case the mode will grow exponentially. The exponential growth continues until the $\gamma_{211}$ becomes almost zero due to black hole spin down in which case the cloud is saturated, or the interaction terms in Eq. (44) become significant. In the former case, the later evolution would be similar to that of free fields. In the later case, the exponential growth could be interrupted by the interaction corrections, and the superradiant cloud reaches a quasiequilibrium stage. By requiring the lhs of Eq. (44) to be zero, one can define the dimensionless occupation number at the quasiequilibrium stage $\epsilon_{eq}$. We expect the cloud reaches the quasi-equilibrium stage before saturation if $\epsilon_{eq}<\epsilon_{max}$.
Although different processes may take place simultaneously as shown in Fig. 7, typically only one process dominates the correction, which allows us to simplify the equation by neglecting the subdominating correction. In Fig. 8, we show the parameter space of the dominating process, with the contours denoting the critical $\tilde{\lambda}_{c}$ defined as $\epsilon_{eq}(\tilde{\lambda}_{c})=\epsilon_{max}$. Namely, the cloud would reach the quasiequilibrium stage before it saturates as in the single-field case if $\tilde{\lambda}>\tilde{\lambda}_{c}$. Here $\tilde{\lambda}$ relates to $\lambda$ via
@@
Given the value of $\tilde{\lambda}_{c}$, we find that the cloud reaches the equilibrium stage with very tiny $\epsilon_{eq}$ even if $\lambda/\mu\ll 1$. Depending on the dominating process, the subsequent evolution may be slightly different, shown in Fig. 10, and will be discussed as follows.
@@ IV.1 Case A
Without the interruption of other superradiant modes, the quasiequilibrium stage continues until $\epsilon_{eq}$ approaches to 0 as the black hole spin approaches to the critical spin ${\tilde{a}}_{crit}$. While Eq. (50) can be solved analytically, we shall not write down the lengthy expression of $a$, but are rather interested in the evolution when $\tilde{a}$} is close to ${\tilde{a}}_{crit}$, in which case
@@ IV.2 Case B
with $\epsilon_{eq}\simeq(\gamma_{211}/\gamma_{R})^{1/2}$. While $\epsilon_{211}$ could also reach $\epsilon_{eq}$ as in case A, the equilibrium may not be always stable. Considering a small deviation $\delta\epsilon$ from $\epsilon_{eq}$ and by perturbing Eq. (53), we have
@@
Comparing the two time scales $\tau_{var}$ and $\tau_{re}$ leads to a threshold in terms of black hole spin
@@
below which the quasiequilibrium stage becomes unstable. If the quasiequilibrium is stable as $a$ approaches $a_{crit}$, the cloud decays with $\epsilon_{211}\propto t^{-1}$. Otherwise, $\epsilon^{\prime}_{211}$ is dominated by the $\gamma_{R}\epsilon_{211}^{3}$ term, and the cloud decays with $\epsilon_{211}\propto t^{-1/2}$.
@@ IV.3 Case C
In region C of Fig. 8, the $\gamma_{B}$ term typically first dominates the interaction correction. The evolution is similar to that in case A, except that the black hole spin stops decreasing once reaching the quasiequilibrium stage. The black hole spin and the occupation number remain almost constant until the $\gamma_{R}$ term becomes significant and takes over the evolution.
@@ IV.4 Cloud collapse
and investigating how $R$ deviates from the Bohr radius $r_{B}$ in the presence of the effective self-interaction. In the nonrelativistic limit, the action of $\bar{\psi}$} reduces to
@@
with $\rho_{\rm BH}=M\delta^{3}({\bf r})$. $\varphi$ is a mediator. Taking into account the contribution of $s$, $t$, and $u$ channels, we have
@@
where $\tilde{R}=R/r_{B}$ and $I_{p}$ is a dimensionless factor containing contribution of $\ell m=22,20,00$, with its numerical values of each component shown in Fig. 11. The four terms in Eq. (61) correspond to kinetic energy, black hole gravity, quartic interaction and cloud self-gravity respectively. The extrema of $V_{\rm eff}$ is located at
@@ V Implications on Observations
Even with a sufficient growth of the cloud, the existence of the other fields can change the evolution of the cloud, including its maximum mass, decay rate and lifetime, and hence leads to different gravitational wave signals compared to the single-field case. These effects should be taken into account in when searching for dark particles with black hole superradiance.
@@ VI Conclusion and Discussion
In this work, we investigated black hole superradiance of interacting multiple fields. Taking a two-field model for demonstration, we discuss the possible interactions between the bound and unbound states of the two fields, and their impacts on the superradiant growth via a perturbative approach. We further analyze the superradiant growth of the fastest mode in the presence of the interactions. We find that the superradiant mode typically experiences three stages: the exponential growth stage, the quasiequilibrium stage and the power-law decay stage. We also find that the superradiance is typically suppressed if the superradiant field couples to another field even with a very weak coupling strength.
While we are taking a two-field model for demonstration, inclusion of more fields should not alter the picture essentially. This is especially the case when there is a large hierarchy in the field mass such that one of the fields grows efficiently, and another field contributes most to the interaction corrections. Nevertheless, one could consider the superradiant growth of the other modes, for example, ${\bar{\psi}}_{322}$, ${\bar{\varphi}}_{211}$ or ${\bar{\varphi}}_{322}$ in our setup, which may play a role in the late evolution of ${\bar{\psi}}_{211}$ in certain parameter regions. Including such modes into the discussion is beyond the scope of this work, and will be studied in future work. It is also interesting to extend the discussion to the spin-1/2 and spin-1 fields, and to consider other possible interactions and their effects on superradiant growth.
It is possible that the dark sector consists of many species of interacting dark particles, rather than just one species of dark particle. As we demonstrated in this work that the existence of the other fields can have important impacts on the observational signatures of superradiance, the superradiance based constraints on dark particles derived from single-field analyses should be revised in the presence of interactions. The observational consequences of multifield black hole superradiance, such as black hole spin distribution, gravitational wave emission from the clouds, and other potential signals of superradiant cloud, are also interesting questions to be studied in future.
@@ Appendix A Ultralight scalar fields in the weak gravity limit of the Kerr black hole background
with $\hat{L}^{2}$ being the total angular momentum operator. In the second line of the above equation, we have used the equation of free $\psi$ at zeroth order of $GM/r$ to eliminate the $\partial_{t}^{2}$ operator. For nonrelativistic states, the wave number $k$ is much less than $\mu$. Therefore, $\hat{K}^{2}\sim k^{2}$ can be neglected, and we have
@@
for nonrelativistic states in the weak gravity limit.
@@
The general solution to Eq. (69) is well known in the literature. By imposing the regular boundary condition, one can select a set of relevant eigenstates $\{u_{n\ell m}\}$ and $\{u_{k\ell m}\}$, which satisfy
@@
In the literature, the bound states are also often described with $\bar{\psi}$} defined by the nonrelativistic ansatz (6). By substituting ansatz (6) into Eq. (67) and keeping leading order terms in $\alpha$, we can find that $\bar{\psi}$} satisfies
@@
with $\xi_{n\ell m}\simeq\omega_{n\ell m}-\mu$. For relativistic states, one should take $\hat{K}^{2}$ into account. Nevertheless, when we calculate the radiation power of the $3{\bar{\psi}}$ process, we take the flat-space approximation for simplicity, in which case the eigenfunction of the unbound states can be obtained by taking $u_{k\ell m}$ to the limit of $kr_{B}\rightarrow\infty$, and the radial function $R_{k\ell}$ reduces to the spherical Bessel function.
Appendix B Green’s function method
In this appendix, we shall show how to solve the inhomogeneous Klein-Gordon equation [^5] Here we use the $\psi$ field to demonstrate the method, while in the main text, the Green’s function method is applied to solve Eq. (17) and Eq. (34) for the $\varphi$ field.
@@ Appendix B Green’s function method
on the Kerr background with the Green’s function method. We shall set the Newton’s constant $G=1$ in this appendix to avoid confusion with the Green’s function. In the Boyer-Lindquist coordinates, the equation is separable with
@@
B.1 The radial Green’s function
The radial function can be solved with the radial Green’s function,
@@ B.1 The radial Green’s function
where $r_{*}$ is the tortoise coordinate. The Green’s function can be constructed as
@@
which describes the ingoing wave at infinity. Since there are only two independent solutions, the above solutions are related by
@@ B.2 Relation to the eigenstates
Figure 13: Analyticity of $\zeta(z)$ and the integral contour in integral (87). The branch cut is chosen at $(-\infty,-\mu]$ and $\mu,\infty)$. The left plot shows the principal Riemann sheet, and the right plot shows the second Riemann sheet. There are poles at $\pm\Omega$ (the red dots), and at the zeros of Wronskian (the black dots). The bound state poles (${\rm Re}\omega<\mu$) are on the principle Riemann sheet, while the qausinormal mode poles are on the second Riemann sheet.
Now we show that Eq. (80) can be expanded with the eigenfunctions of the homogeneous radial equation, i.e., Eq. (B) with $I=0$. We shall show this by constructing an integral
@@
Recall that $\Omega$ is the forced oscillation frequency of the source term $s(x)$. The integration contour and the analyticity of $\zeta(z)$ are shown in Fig. 13. Because of the multivalue of $k=\pm\sqrt{z^{2}-\mu^{2}}$, $\zeta(z)$ has two branch points $z=\pm\mu$ and we define the integral in the Riemann sheet of ${\rm Im}\,k>0$ to fit the outgoing boundary condition, which means the branch cut is $(-\infty,-\mu]$ and $\mu,\infty)$.
@@
such that it approaches the normalized eigenfunctions discussed in Appendix. A in the weak gravity limit.
@@
For $s(x)=\tilde{s}(\mathbf{r})e^{-i\Omega t}+c.c.$ with general angular dependence, one can consider the full Green’s function [62]
@@
and use Eq. (93) when computing the radial part. Specifically, in the weak gravity limit, we have $A_{\mathrm{in}}A_{\mathrm{out}}=1$, $f_{n}(r)S_{\omega\ell m}(\theta)e^{im\phi}$ and $f_{H}(r)S_{\omega\ell m}(\theta)e^{im\phi}$ reduce to $u_{n\ell m}(\mathbf{r})$ and $u_{k\ell m}(\mathbf{r})$ discussed in Appendix. A respectively. Therefore, the solution to Eq. (19) can be given by Eq. (LABEL:eq:expPhi).